# Straggler-Resilient Decentralized Learning via Adaptive Asynchronous Updates

### v1

@@ C.3 Auxiliary Lemmas
Lemma 2 (Lemma 4 in [nedic2009distributed]).
@@
$\left|\frac{1}{N}-\Phi_{k,s}(i,j)\right|\leq 2\frac{(1+\beta^{-NB})}{1-\beta^{NB}}(1-\beta^{NB})^{(k-s)/NB},$ (21)
@@
C.4 Assumption on Adaptive Asynchronous Updates
@@
We note that we focus on stragglers caused by heterogeneous computation capabilities among workers as in [ananthanarayanan2013effective, ho2013more, lian2017can, lian2018asynchronous, luo2020prague], while the communication bandwidth is not a bottleneck. Hence the major time consumption comes from the calculation of gradients. First, we prove that DSGD-AAU achieves a linear speedup for convergence in terms of iterations. In such case, the communication does not affect the result with respect to the convergence iteration. However, the communication time will impact the wall-clock time for convergence. To the best of knowledge, Prague [luo2020prague] is perhaps the first work to separately measure the computation time and communication time by running different number of workers and found that the communication time is much smaller and dominated by the computation time. We further explore this and estimate the communication time in our system given the network speed. For instance, in our system, the network speed is about 20GB/s given the hardware of our server (memory card), and the communication time only accounts for 0.14%-4% of the total time measured in communication and computation time under the CIFAR-10 dataset. Moreover, our convergence time in the numerical evaluation considers the total time of the whole process, which includes both computation and communication time. Exactly same thing has been done in Prague [luo2020prague], AD-PSGD [lian2018asynchronous], and Li et al. [li2020taming, li2020breaking].
Appendix D Additional Experimental Results
@@
In this section, we provide the experiment details, parameter settings and some additional experimental results of the experiment setting in Section 5.
@@
Dataset and Model. We evaluate the performance of different algorithms on both the classification tasks and next-character prediction tasks. For the classification tasks, we use CIFAR-10 [krizhevsky2009learning], MNIST [lecun1998gradient] and Tiny-ImageNet [le2015tiny, russakovsky2015imagenet] as the evaluation datasets for both i.i.d. and non-i.i.d. versions. However, we mainly focus on the non-i.i.d. versions since data are generated locally at the workers based on their circumstances in real-world systems (see Section 3). For the task of next-character prediction, we use the dataset of The Complete Works of William Shakespeare [mcmahan2017communication] (Shakespeare).
@@
The CIFAR-10 dataset consists of $60,000$ $32\times 32$ color images in $10$ classes where $50,000$ samples are for training and the other $10,000$ samples for testing. The MNIST dataset contains handwritten digits with $60,000$ samples for training and $10,000$ samples for testing. The Tiny-ImageNet [le2015tiny] has a total of $200$ classes with $500$ images each class for training and $50$ for testing, which is drawn from the ImageNet dataset [russakovsky2015imagenet] and down-sampled to $64\times 64$ For MNIST, we split the data based on the digits they contain in the dataset, i.e., we distribute the data to all workers such that each worker contains only a certain class of digits with the same number of training/testing samples (e.g., same ideas were used in Yang et al. [yang2021achieving]). We also use the idea suggested by McMahan et al. [mcmahan2017communication] and Yang et al. [yang2021achieving] to deal with CIFAR-10 dataset. We use $N=128$ workers as an example. We sort the data by label of different classes (10 classes in total), and split each class into $N/2=64$ pieces. Each worker will randomly select 5 classes of data partitions to forms its local dataset.
For the classification tasks, we consider four representative models: a fully-connected neural network with 2 hidden layers (2NN), ResNet-18 [he2016deep], VGG-13 [simonyan2015very] and AlexNet [krizhevsky2012imagenet], as Table 3, Table 4, Table 5 and Table 6, respectively. For the next-character prediction task, we use an LSTM language model as Table 7 as in [kim2016character].
Parameter Shape Layer hyper-parameter
layer1.fc1.weight $3072\times 256$ N/A
layer1.fc1.bias 256 N/A
layer2.fc2.weight $256\times 256$ N/A
layer2.fc2.bias 256 N/A
layer3.fc3.weight $256\times 10$ N/A
layer3.fc3.bias 10 N/A
Table 3: Detailed information of the 2NN architecture used in our experiments. All non-linear activation function in this architecture is ReLU. The shapes for convolution layers follow $(C_{in},C_{out},c,c)$.
Parameter Shape Layer hyper-parameter
layer1.conv1.weight $3\times 3,64$ stride:1; padding: 1
layer1.conv1.bias 64 N/A
batchnorm2d 64 N/A
layer2.conv2 $\begin{bmatrix}3\times 3,&64\\ 3\times 3,&64\\ \end{bmatrix}\times 2$ stride:1; padding: 1
layer3.conv3 $\begin{bmatrix}3\times 3,&128\\ 3\times 3,&128\\ \end{bmatrix}\times 2$ stride:1; padding: 1
layer4.conv4 $\begin{bmatrix}3\times 3,&256\\ 3\times 3,&256\\ \end{bmatrix}\times 2$ stride:1; padding: 1
layer5.conv5 $\begin{bmatrix}3\times 3,&512\\ 3\times 3,&512\\ \end{bmatrix}\times 2$ stride:1; padding: 1
pooling.avg N/A N/A
layer6.fc6.weight $512\times 10$ N/A
layer6.fc6.bias 10 N/A
Table 4: Detailed information of the ResNet-18 architecture used in our experiments. All non-linear activation function in this architecture is ReLU. The shapes for convolution layers follow $(C_{in},C_{out},c,c)$.
Parameter Shape Layer hyper-parameter
layer1.conv1.weight $3\times 64\times 3\times 3$ stride:1; padding: 1
layer1.conv1.bias 64 N/A
pooling.max N/A kernel size:2; stride: 2
layer2.conv2.weight $32\times 64\times 3\times 3$ stride:1; padding: 1
layer2.conv2.bias 64 N/A
pooling.max N/A kernel size:2; stride: 2
layer3.conv3.weight $64\times 128\times 3\times 3$ stride:1; padding: 1
layer3.conv3.bias 128 N/A
layer4.conv4.weight $128\times 128\times 3\times 3$ stride:1; padding: 1
layer4.conv4.bias 128 N/A
pooling.max N/A kernel size:2; stride: 2
layer5.conv5.weight $128\times 256\times 3\times 3$ stride:1; padding: 1
layer5.conv5.bias 256 N/A
layer6.conv6.weight $256\times 256\times 3\times 3$ stride:1; padding: 1
layer6.conv6.bias 256 N/A
pooling.max N/A kernel size:2; stride: 2
layer7.conv7.weight $256\times 512\times 3\times 3$ stride:1; padding: 1
layer7.conv7.bias 512 N/A
layer8.conv8.weight $512\times 512\times 3\times 3$ stride:1; padding: 1
layer8.conv8.bias 512 N/A
pooling.max N/A kernel size:2; stride: 2
layer9.conv9.weight $512\times 512\times 3\times 3$ stride:1; padding: 1
layer9.conv9.bias 512 N/A
layer10.conv10.weight $512\times 512\times 3\times 3$ stride:1; padding: 1
layer10.conv10.bias 512 N/A
pooling.max N/A kernel size:2; stride: 2
dropout N/A p=20%
layer11.fc11.weight $4096\times 512$ N/A
layer11.fc11.bias 512 N/A
layer12.fc12.weight $512\times 512$ N/A
layer12.fc12.bias 512 N/A
dropout N/A p=20%
layer13.fc13.weight $512\times 10$ N/A
layer13.fc13.bias 10 N/A
Table 5: Detailed information of the VGG-13 architecture used in our experiments. All non-linear activation function in this architecture is ReLU. The shapes for convolution layers follow $(C_{in},C_{out},c,c)$.
Parameter Shape Layer hyper-parameter
layer1.conv1.weight $3\times 64\times 3\times 3$ stride:2; padding: 1
layer1.conv1.bias 64 N/A
pooling.max N/A kernel size:2; stride: 2
layer2.conv2.weight $64\times 192\times 3\times 3$ stride:1; padding: 1
layer2.conv2.bias 64 N/A
pooling.max N/A kernel size:2; stride: 2
layer3.conv3.weight $192\times 384\times 3\times 3$ stride:1; padding: 1
layer3.conv3.bias 128 N/A
layer4.conv4.weight $384\times 256\times 3\times 3$ stride:1; padding: 1
layer4.conv4.bias 128 N/A
layer5.conv5.weight $256\times 256\times 3\times 3$ stride:1; padding: 1
layer5.conv5.bias 256 N/A
pooling.max N/A kernel size:2; stride: 2
dropout N/A p=20%
layer6.fc6.weight $1024\times 4096$ N/A
layer6.fc6.bias 512 N/A
dropout N/A p=20%
layer7.fc7.weight $4096\times 4096$ N/A
layer7.fc7.bias 512 N/A
layer8.fc8.weight $4096\times 10$ N/A
layer8.fc8.bias 10 N/A
Table 6: Detailed information of the AlexNet architecture used in our experiments. All non-linear activation function in this architecture is ReLU. The shapes for convolution layers follow $(C_{in},C_{out},c,c)$.
Parameter Shape Layer hyper-parameter
layer1.embeding $80\times 256$ N/A
layer2.lstm $256\times 512$ num_layers=2, batch_first=True
dropout N/A p=5%
layer3.fc.weight $512\times 80$ N/A
layer3.fc.bias 80 N/A
Table 7: Detailed information of the LSTM model used in our experiments.
@@
Our experiments are running using the following setups. Model: AMD TRX40 Motherboard. System: Ubuntu 20.04. CPU: AMD Ryzen threadripper 3960x 24-core processor x 48. GPU: NVIDIA RTX A6000. Network Speed: 20GB/s. Language: Python3, Tensorflow 2.0. In real systems, there are several reasons leading to the existing of stragglers, e.g., heterogeneous hardware, hardware failure, imbalanced data distributions among tasks and different OS effects, resource contention, etc. Since we run our experiments of different number of workers in one sever as in many other literatures, e.g., Li et al. [li2020taming, li2020breaking] and motivated by the idea of AD-PSGD [lian2018asynchronous] and Cipar et al. [cipar2013solving], we randomly select workers as stragglers in each iteration. If a worker is selected to be a straggler, then it will sleep for some time in the iteration (e.g., the sleep time could be 6x of the average one local computation time). We also investigate the impact of the sleep time and the percentage of stragglers in our ablation study below. Such techniques have been widely used in the literature such as AD-PSGD [lian2018asynchronous], Prague [luo2020prague] and Li et al. [li2020taming, li2020breaking].
@@
Table 8: The test accuracy of different models using different non-i.i.d. datasets with 128 workers.
@@
Table 9: The test accuracy of ResNet-18 on non-i.i.d. CIFAR-10, MNIST, and Tiny-ImageNet when trained for 50 seconds, 10 seconds and 30 seconds, respectively; and the test accuracy of LSTM on non-i.i.d. Shakespeare when trained for 30 seconds.
Test Accuracy. Complementary to Table 1 in the main paper, we summarize the test accuracy of different models using different non-i.i.d. datasets with 128 workers in Table 8. Again, we observe that DSGD-AAU consistently outperforms all state-of-the-art baselines in consideration.
Complementary to Table 2 in the main paper, we summarize the test accuracy of different models over different non-i.i.d. datasets when trained for a limited time in Table 9 with different number of workers. For example, for the image classification tasks, we train ResNet-18 for 50 seconds on non-i.i.d. CIFAR-10, 10 seconds on non-i.i.d. MNIST, and 50 seconds for non-i.i.d. Tiny-ImageNet. For the next-character prediction task, we train LSTM on Shakespeare for 30 seconds. Again, we observe that for a limited training time, DSGD-AAU consistently achieves a higher test accuracy compared to AGP, AD-PSGD and Prague.
Complementary to Tables 8 and 9, the corresponding results when trained on the i.i.d. version of datasets are presented in Tables 10 and 11, respectively. Similar observations can be made and hence we omit the discussions here.
@@
Table 10: The test accuracy of different models using different i.i.d. datasets with 128 workers.
@@
Table 11: The test accuracy of ResNet-18 on i.i.d. CIFAR-10, MNIST, and Tiny-ImageNet when trained for 50 seconds, 10 seconds and 30 seconds, respectively; and the test accuracy of LSTM on Shakespeare when trained for 30 seconds.
Speedup and Communication. Complementary to Figure 5 in the main paper, we present the corresponding results for 2-NN, AlexNet, and VGG-13 in Figures 67 and 8, respectively. Again, we observe that DSGD-AAU consistently outperforms baseline methods and achieves the best speedup at no cost of additional communication.
Figure 6: Speedup and communication for 2-NN on non-i.i.d. CIFAR-10 with different number of workers.
Figure 7: Speedup and communication for AlexNet on non-i.i.d. CIFAR-10 with different number of workers.
Figure 8: Speedup and communication for VGG-13 on non-i.i.d. CIFAR-10 with different number of workers.
@@
Similar observations can be made when consider the i.i.d. CIFAR-10, which is shown in Figures 11 and 12, as well as other models and datasets. Hence we omit the results and discussions here.
Figure 9: Ablation study for VGG-13 on non-i.i.d. CIFAR-10.
Figure 10: Ablation study for VGG-13 when trained for 50 seconds on non-i.i.d. CIFAR-10.
Figure 11: Ablation study for VGG-13 on i.i.d. CIFAR-10.
Figure 12: Ablation study for VGG-13 when trained for 50 seconds on i.i.d. CIFAR-10.

### v2

@@ B.3. Auxiliary Lemmas
Lemma B.3 (Lemma 4 in (Nedic and Ozdaglar, 2009)).
@@
(21) $\left|\frac{1}{N}-\Phi_{k,s}(i,j)\right|\leq 2\frac{(1+\beta^{-NB})}{1-\beta^{NB}}(1-\beta^{NB})^{(k-s)/NB},$
@@
B.4. Assumption on Adaptive Asynchronous Updates
@@
We focus on stragglers caused by heterogeneous computation capabilities among workers as in (Ananthanarayanan et al., 2013; Ho et al., 2013; Lian et al., 2017, 2018; Luo et al., 2020), while the communication bandwidth is not a bottleneck. Hence the major time consumption comes from the calculation of gradients. First, we prove that DSGD-AAU achieves a linear speedup for convergence in terms of iterations. In such case, the communication does not affect the result with respect to the convergence iteration. However, the communication time will impact the wall-clock time for convergence. To the best of knowledge, Prague (Luo et al., 2020) is perhaps the first work to separately measure the computation time and communication time by running different number of workers and found that the communication time is much smaller and dominated by the computation time. We further explore this and estimate the communication time in our system given the network speed. For instance, in our system, the network speed is about 20GB/s given the hardware of our server (memory card), and the communication time only accounts for 0.14%-4% of the total time measured in communication and computation time under the CIFAR-10 dataset. Moreover, our convergence time in the numerical evaluation considers the total time of the whole process, which includes both computation and communication time. Exactly same thing has been done in Prague (Luo et al., 2020), AD-PSGD (Lian et al., 2018), and Li et al. (Li et al., 2020a, b).
Appendix C Additional Experimental Results
@@
We provide the experiment details, parameter settings and some additional experimental results of the experiment setting in Section 6.
@@
Dataset and Model. We evaluate the performance of different algorithms on both the classification tasks and next-character prediction tasks. For the classification tasks, we use CIFAR-10 (Krizhevsky et al., 2009), MNIST (LeCun et al., 1998) and Tiny-ImageNet (Le and Yang, 2015; Russakovsky et al., 2015) as the evaluation datasets for both i.i.d. and non-i.i.d. versions. However, we mainly focus on the non-i.i.d. versions since data are generated locally at the workers based on their circumstances in real-world systems (see Section 2). For the task of next-character prediction, we use the dataset of The Complete Works of William Shakespeare (McMahan et al., 2017) (Shakespeare).
@@
The CIFAR-10 dataset consists of $60,000$ $32\times 32$ color images in $10$ classes where $50,000$ samples are for training and the other $10,000$ samples for testing. The MNIST dataset contains handwritten digits with $60,000$ samples for training and $10,000$ samples for testing. The Tiny-ImageNet (Le and Yang, 2015) has a total of $200$ classes with $500$ images each class for training and $50$ for testing, which is drawn from the ImageNet dataset (Russakovsky et al., 2015) and down-sampled to $64\times 64$ For MNIST, we split the data based on the digits they contain in the dataset, i.e., we distribute the data to all workers such that each worker contains only a certain class of digits with the same number of training/testing samples (Yang et al., 2021). We also use the idea suggested by McMahan et al. (McMahan et al., 2017) and Yang et al. (Yang et al., 2021) to deal with CIFAR-10 dataset. We use $N=128$ workers as an example. We sort the data by label of different classes (10 classes in total), and split each class into $N/2=64$ pieces. Each worker will randomly select 5 classes of data partitions to forms its local dataset.
@@
Our experiments are running using the following setups. Model: AMD TRX40 Motherboard. System: Ubuntu 20.04. CPU: AMD Ryzen threadripper 3960x 24-core processor x 48. GPU: NVIDIA RTX A6000. Network Speed: 20GB/s. Language: Python3, Tensorflow 2.0. In real systems, there are several reasons leading to the existing of stragglers, e.g., heterogeneous hardware, hardware failure, imbalanced data distributions among tasks and different OS effects, resource contention, etc. Since we run our experiments of different number of workers in one sever as in many other literatures, e.g., Li et al. (Li et al., 2020a, b) and motivated by the idea of AD-PSGD (Lian et al., 2018) and Cipar et al. (Cipar et al., 2013), we randomly select workers as stragglers in each iteration. If a worker is selected to be a straggler, then it will sleep for some time in the iteration (e.g., the sleep time could be 6x of the average one local computation time). We also investigate the impact of the sleep time and the percentage of stragglers in our ablation study below. Such techniques have been widely used in the literature such as AD-PSGD (Lian et al., 2018), Prague (Luo et al., 2020) and Li et al. (Li et al., 2020a, b).
@@
Table 3. The test accuracy of different models using different non-i.i.d. datasets with 128 workers.
@@
Table 4. The test accuracy of ResNet-18 on non-i.i.d. CIFAR-10, MNIST, and Tiny-ImageNet trained for 50 seconds, 10 seconds and 30 seconds, respectively; and the test accuracy of LSTM on non-i.i.d. Shakespeare trained for 30 seconds.
Test Accuracy. Complementary to Table 1, we summarize the test accuracy of different models using different non-i.i.d. datasets with 128 workers in Table 3. Again, we observe that DSGD-AAU consistently outperforms all state-of-the-art baselines in consideration. Complementary to Table 2, we summarize the test accuracy of different models over different non-i.i.d. datasets when trained for a limited time in Table 4 with different number of workers. For example, for the image classification tasks, we train ResNet-18 for 50 seconds on non-i.i.d. CIFAR-10, 10 seconds on non-i.i.d. MNIST, and 50 seconds for non-i.i.d. Tiny-ImageNet. For the next-character prediction task, we train LSTM on Shakespeare for 30 seconds. Again, we observe that for a limited training time, DSGD-AAU consistently achieves a higher test accuracy compared to AGP, AD-PSGD and Prague. Complementary to Tables 3 and 4, the corresponding results when trained on the i.i.d. version of datasets are presented in Tables 5 and 6, respectively. Similar observations can be made.
@@
Table 5. The test accuracy of different models using different i.i.d. datasets with 128 workers.
@@
Table 6. The test accuracy of ResNet-18 on i.i.d. CIFAR-10, MNIST, and Tiny-ImageNet when trained for 50 seconds, 10 seconds and 30 seconds, respectively; and the test accuracy of LSTM on Shakespeare when trained for 30 seconds.
Speedup and Communication. Complementary to Figure 5, we present corresponding results for 2-NN, AlexNet, and VGG-13 in Figures 67 and 8. Again, we observe that DSGD-AAU consistently outperforms baseline methods and achieves the best speedup at no cost of additional communication.
Figure 6. Speedup and communication for 2-NN on non-i.i.d. CIFAR-10 with different number of workers.
Figure 7. Speedup and communication for AlexNet on non-i.i.d. CIFAR-10 with different number of workers.
Figure 8. Speedup and communication for VGG-13 on non-i.i.d. CIFAR-10 with different number of workers.
@@
Similar observations can be made when consider the i.i.d. CIFAR-10, as well as other models and datasets. Hence we omit the results and discussions here.
Figure 9. Ablation study for VGG-13 on non-i.i.d. CIFAR-10.
Figure 10. Ablation study for VGG-13 when trained for 50 seconds on non-i.i.d. CIFAR-10.